Question 5: what is the new EOQ if a product whose EOQ is 500 experiences a 30% increase in demand
Answer #1

Economic Order Quantity (EOQ) is the ideal quantity a company needs to order for its inventory to reduce invariable costs such as holding charges, storage charges, cost to order etc

The formula for EOQ is as follows= \sqrt{\frac{2DS}{H}} where D is demand, S is the cost of order and H is the cost of holding

Since units of EOQ is directly proportional to the square root of demand i.e. D, an increase in 30% of demand means existing EOQ units increases by.\sqrt{\frac{2(1.3D))S}{H}}

Steps:-

Step 1:- EOQ Units= \sqrt{\frac{2DS}{H}}

Step 2:- Since demand increases by 30%, therefore new demand is 1.3D. Substituting the new value in the equation

New EOQ Units= \sqrt{\frac{2(1.3D))S}{H}}

Step 3:- New EOQ Units= \sqrt{1.3}*\sqrt{\frac{2DS}{H}}

Substituting the value of 500 units for earlier EOQ Units, new EOQ units will be \sqrt{1.3}*500

Therefore, New EOQ Units= 570.08i.e. 570 Units